# AdaWorld: Learning Adaptable World Models with Latent Actions

### v1

@@ 3.2.1 Simulation Quality
Setup. We choose three action-controlled environments (i.e., Habitat (Savva et al., 2019), Minecraft, DMLab (Beattie et al., 2016)) that are not included in our training dataset. To demonstrate the adaptability with restricted labels, we collect only 100 samples for each action in every environment. Using the limited interaction data, we then finetune all compared world models for 800 steps with a batch size of 32 and a learning rate of $5\times 10^{-5}$. The learning rate for the pretrained weights is discounted by a factor of 0.1. We randomly initialize action embeddings for the action-agnostic pretraining baseline. For the other three models, we use the averaged action conditions extracted from the 100 samples to initialize the action embeddings as described in Sec. 2.3. The adaptation quality is evaluated via PSNR (Hore & Ziou, 2010) and LPIPS (Zhang et al., 2018).
@@
To further demonstrate our sample efficiency and finetuning efficiency, we select Minecraft as the environment and conduct more comparative experiments using different sample numbers and finetuning steps. We show the evolving curves of PSNR in Figure 6. In all cases, AdaWorld performs much better at the beginning and improves significantly faster after a few finetuning steps. This suggests that our approach provides a superior initialization for efficient world model adaptation compared to conventional pretraining methods.
3.2.2 Visual Planning
Table 3: Visual planning results. For all selected scenes from four Procgen environments, we report the success rate and standard error with 5 random seeds. AdaWorld achieves higher average success rates than other baselines as well as Q-learning even without training. Oracle: We use the ground truth environments for MPC, which indicate the upper bound of this planning strategy.
@@ 3.2.2 Visual Planning
   w/o finetune 38.67$\pm$0.8268.00$\pm$0.9241.33$\pm$1.1131.33$\pm$1.0244.83$\pm$0.28%
   w/ finetune 66.67$\pm$1.6758.67$\pm$1.0268.00$\pm$0.6933.33$\pm$1.5556.67$\pm$0.44%
Q-learning 22.67$\pm$1.5847.33$\pm$2.744.67$\pm$0.3334.00$\pm$2.5227.17$\pm$0.26%
Oracle (GT env.) 86.67$\pm$1.2977.33$\pm$1.0984.67$\pm$1.1974.00$\pm$1.6380.67$\pm$0.43%
Setup. After learning action controls, world models can be utilized for planning. To evaluate the impact of adaptability on planning performance, we compare AdaWorld with the common action-agnostic pretraining using a sampling-based model predictive control (MPC) optimized by Cross-Entropy Method (De Boer et al., 2005; Chua et al., 2018). The MPC planning and optimization procedure are deferred to Sec. B.4. We define a goal-reaching task based on the Procgen benchmark (Cobbe et al., 2020) and select 30 scenes from each of four environments (Heist, Jumper, Maze, CaveFlyer). This ensures that the specified goals can be reached within an acceptable number of steps (more details in Appendix E). For each scene, we randomly collect 100 samples for each action in the default action space (LEFT, DOWN, UP, RIGHT). Based on the collected samples, the pretrained world models are finetuned for 500 steps with a batch size of 32 and a learning rate of $5\times 10^{-5}$. We then use the finetuned world models to perform MPC planning in the selected scenes. The reward is defined as the cosine similarity between the predicted observations and the image of the final state. The MPC process is deemed successful if the agent reaches the final state within 20 steps.
Results. We present the success rates averaged over 5 random seeds in Table 3. While the action-agnostic baseline performs similarly to random planning, AdaWorld substantially increase the success rates across all environments. This indicates that our approach not only adapts more efficiently but also enhances visual planning performance.
@@
3.3 Ablation Study
Table 4: Impacts of training data diversity. Increasing data diversity enhances generalization to new domains.
@@ 3.3 Ablation Study
Averaged action embedding. We ablate the averaged action embedding proposed in 2.3 with random initialization and adapt the resultant model to the unseen Minecraft environment. From Figure 6, we observe that randomly initializing the action embeddings of AdaWorld results in a quality drop at the beginning. Nevertheless, it is noteworthy that although our variant is slightly worse than the action-agnostic pretraining baseline when the finetuning begins, it rapidly surpasses the action-agnostic baseline after just 200 steps. This is because AdaWorld has learned a highly adaptable control interface through action-aware pretraining, allowing it to efficiently adapt to an unseen environment by simply fitting the action embeddings for the new action space.
@@
4 Related Work
4.1 World Models
Driven by advances in deep generative models (Peebles & Xie, 2023; Yang et al., 2025), world models have made encouraging strides in recent years. They aim to replicate the transitions of the world through generation (Sutton, 1991; Ha & Schmidhuber, 2018; Yang et al., 2024c; Lu et al., 2024, 2025; Agarwal et al., 2025), enabling agents to perform decision making entirely in imagination even when confronted with unseen situations (Kaiser et al., 2020; Kim et al., 2020; Micheli et al., 2023; Hafner et al., 2023; Du et al., 2023; Mazzaglia et al., 2024). This capability has proven beneficial for gaming agents (Alonso et al., 2024; Pearce et al., 2024; He et al., 2025), autonomous vehicles (Hu et al., 2023; Yang et al., 2024a; Gao et al., 2024; Wang et al., 2024; Bar et al., 2025; Hassan et al., 2025), and real robots (Seo et al., 2023; Wu et al., 2022; Ko et al., 2024; Du et al., 2024; Zhen et al., 2024; Zhou et al., 2024b; Wu et al., 2024; Zhu et al., 2024; Zhou et al., 2024a; Bu et al., 2024; Wang et al., 2025; Qi et al., 2025; Wu et al., 2025).
4.2 Latent Action from Videos
Humans can comprehend the notion of various actions by observing others to behave (Rizzolatti et al., 1996; Rybkin et al., 2019; Schmeckpeper et al., 2020). To develop intelligent agents, it is intriguing to automatically extract similar action primitives from videos. Some approaches align human and robot manipulation videos in a shared space to derive meaningful action prototypes (Xu et al., 2023b). However, they require collecting semantically paired videos, which limits their scalability for most real-world tasks.
To avoid reliance on video pairs, some works propose extracting implicit latent actions through fully unsupervised learning (Edwards et al., 2019; Menapace et al., 2021, 2022; Ye et al., 2023; Schmidt & Jiang, 2024; Sun et al., 2024; Villar-Corrales & Behnke, 2025). Nevertheless, they mainly focus on a single environment with limited complexity. While more recent works extend similar approaches to multiple datasets (Cui et al., 2024; Chen et al., 2024b, c; Ye et al., 2025; Bu et al., 2025; Ren et al., 2025), they mainly leverage latent actions as the objectives for behavior cloning. Consequently, they always adopt discrete actions to fit the policy networks (Kim et al., 2024), leading to significant ambiguity due to discretization (Nikulin et al., 2025). As a result, the potential of latent actions for adaptable world modeling remains underexplored.
@@ A.2 Data Mixture
Table 5: Data organization. Data sources, estimated frame counts, and mixture ratios for our training.
Category Data Source # Frames Ratios
@@
2D Video Game Gym Retro (Nichol et al., 2018) 1000M 49%
Procgen Benchmark (Cobbe et al., 2020) 144M 2%
@@ B.3 Sampling
B.4 Visual Planning with Model Predictive Control
@@ C.1 Action Transfer
Visualizations. We attach more action transfer results in a number of environments in Figures 11, 12, 13, 14 and 15. Each sample is generated by transferring a latent action sequence with a length of 20.
@@ Appendix E Selected Scenes for Visual Planning
To guarantee an effective evaluation, we conduct an exhaustive search using the ground truth environments to identify viable Procgen test scenes that can be completed in an acceptable number of steps by goal image matching. This results in a total of 120 scenes for our visual planning experiments in Sec. 3.2.2. In Figures 19, 16, 17 and 18, we illustrate the initial states of all selected scenes. The scenes from survival-oriented environments (e.g., BigFish and BossFight) are not considered in our evaluation.

### v2

@@ 3.2.1 Simulation Quality
Setup. To evaluate the simulation quality after adaptation, we choose three environments with discrete action space (Habitat (Savva et al., 2019), Minecraft, DMLab (Beattie et al., 2016)) and one environment with continuous action space (nuScenes (Caesar et al., 2020)) that are not included in our training dataset. Each environment has a validation set consisting of 300 samples, which is used to evaluate the adaption quality in terms of PSNR (Hore & Ziou, 2010) and LPIPS (Zhang et al., 2018). To demonstrate the adaptability with restricted labels, we collect only 100 samples for each action in every discrete environment and 100 trajectories for nuScenes. Using the limited interaction data, we then finetune all compared world models for 800 steps with a batch size of 32 and a learning rate of $5\times 10^{-5}$. The learning rate for the pretrained weights is discounted by a factor of 0.1. For the action-agnostic baseline, we initialize action embeddings with random parameters. For the other three models, we use the averaged action conditions extracted from the 100 samples to initialize the action embeddings as described in Sec. 2.3. Note that for nuScenes, we add a two-layer MLP to map continuous displacements to the latent action interface. The MLP is fine-tuned with limited action-latent action pairs for 3K steps, which takes less than 30 seconds on a single GPU.
@@
To further demonstrate our sample efficiency and finetuning efficiency, we conduct more comparative experiments using different sample numbers and finetuning steps on Minecraft and nuScenes. We show the evolving curves of PSNR in Figure 6. In all cases, AdaWorld performs much better at the beginning and improves significantly faster after a few finetuning steps. This suggests that our approach provides a superior initialization for efficient world model adaptation compared to conventional pretraining methods.
3.2.2 Visual Planning in Games
Table 3: Visual planning results in games. For all selected scenes from four Procgen environments, we report the success rate and standard error with 5 random seeds. AdaWorld achieves higher average success rates than other baselines as well as Q-learning even without finetuning. Oracle: We use the ground truth simulator for MPC, which indicate the upper bound of this planning strategy.
@@ 3.2.2 Visual Planning in Games
   w/o finetune 38.67$\pm$2.0168.00$\pm$2.2541.33$\pm$2.7231.33$\pm$2.5044.83$\pm$1.37%
   w/ finetune 66.67$\pm$4.0958.67$\pm$2.5068.00$\pm$1.6933.33$\pm$3.8056.67$\pm$2.16%
Q-learning 22.67$\pm$3.8747.33$\pm$6.714.67$\pm$0.8134.00$\pm$6.1727.17$\pm$1.27%
Oracle (GT env.) 86.67$\pm$3.1677.33$\pm$2.6784.67$\pm$2.9174.00$\pm$3.9980.67$\pm$2.11%
Table 4: Visual planning results in robot tasks. The success rates and standard errors are obtained over 4 runs for each task from VP2. We also report the aggregated success rates normalized by the scores of the ground truth simulator on the right.
[…]
Robosuite push Open slide Blue button Green button Red button Upright block Aggregate
Act-agnostic 17.50$\pm$0.501.67$\pm$1.675.00$\pm$1.673.33$\pm$0.000.00$\pm$0.001.67$\pm$1.675.03
AdaWorld 63.50$\pm$1.715.83$\pm$2.8529.17$\pm$2.5010.83$\pm$2.5010.00$\pm$2.365.00$\pm$0.9621.54
Setup. After learning action controls, world models can be utilized for planning. To demonstrate the superiority of AdaWorld in planning performance, we first compare it with the action-agnostic baseline in video game environments using sampling-based model predictive control (MPC) optimized by Cross-Entropy Method (De Boer et al., 2005; Chua et al., 2018). The MPC planning and optimization procedure are deferred to Sec. B.4. We define a goal-reaching task based on the Procgen benchmark (Cobbe et al., 2020) and select 30 scenes from each of four environments (Heist, Jumper, Maze, CaveFlyer). This ensures that the specified goals can be reached within an acceptable number of steps (more details in Appendix E). For each scene, we randomly collect 100 samples for each action in the default action space (LEFT, DOWN, UP, RIGHT). Based on the collected samples, the pretrained world models are finetuned for 500 steps with a batch size of 32 and a learning rate of $5\times 10^{-5}$. We then use the finetuned world models to perform MPC planning in the selected scenes. The reward is defined as the cosine similarity between the predicted observations and the image of the final state. The planning is deemed successful if the agent reaches the final state within 20 steps.
Results. Table 3 presents the success rates averaged over 5 random seeds. While the action-agnostic baseline performs similarly to random planning, AdaWorld substantially increase the success rates across all environments. This indicates that our approach not only adapts more efficiently but also enables more effective planning. Visit our project page to see planning demonstrations of agents in games.
@@
3.2.3 Visual Planning in Robot Tasks
Setup. To verify our efficacy in robot control tasks, we pretrain a low-resolution variant of AdaWorld and evaluate the planning performance on the VP2 benchmark (Tian et al., 2023) after adaptation. The planning is also sampling-based and is performed using the model-predictive path integral (MPPI) (Williams et al., 2016; Nagabandi et al., 2020). We focus on a similar compute-efficient setting and finetune the pretrained variant and an action-agnostic baseline for 1K steps. The evaluation is conducted on 100 tabletop Robosuite tasks (Zhu et al., 2020) and 7 RoboDesk tasks (Kannan et al., 2021). More details are in Sec. B.5.
Results. Table 4 reports the success rates on VP2. We omit the Flat block and Open drawer tasks from RoboDesk, as they do not yield meaningful scores under our constrained adaptation setting. The results show that AdaWorld adapts more efficiently with limited finetuning steps and improves the planning performance by a clear margin.
3.3 Ablation and Analysis
Interface initialization. We ablate the latent action initialization approaches in Sec. 2.3 with random initialization and adapt the resultant model to the unseen Minecraft and nuScenes. Figure 6 shows that randomly initializing the control interface of AdaWorld results in a quality drop at the beginning. Nevertheless, it is noteworthy that although our variant is slightly worse than the action-agnostic pretraining baseline when the finetuning begins, it rapidly surpasses the action-agnostic baseline after just 200 steps. This is because AdaWorld has learned a highly adaptable control interface through action-aware pretraining, allowing it to efficiently adapt to an unseen environment by simply fitting the action embeddings for the new action space.
Table 5: Impacts of training data diversity. Increasing data diversity enhances the generalization of latent actions to new domains.
@@ 3.3 Ablation and Analysis
Table 6: Generality of AdaWorld. Applying action-aware pretraining to iVideoGPT also significantly improves its adaptability.
Model BAIR
[…]
iVideoGPT 16.59 0.220
iVideoGPT+AdaWorld 17.40 0.204
@@
Method generality. To demonstrate the generality of our method, we use iVideoGPT (Wu et al., 2024) as a state-of-the-art baseline. iVideoGPT is an action-controlled world model with an autoregressive Transformer architecture. It is pretrained by action-agnostic video prediction and adds a linear projection to learn action control during finetuning. For fair comparison, we implement a variant by conditioning iVideoGPT with our latent actions during pretraining. The training details can be found in Sec. B.6. After finetuning, we compare action-controlled simulation quality on BAIR robot pushing dataset (Ebert et al., 2017) in Table 6. The proposed action-aware pretraining significantly enhances the adaptability of iVideoGPT, suggesting that our method is generally applicable to different world models.
Hyperparameter choice. In Eq. (2), the hyperparameter $\beta$ is adjusted to achieve a good trade-off between expressiveness and context disentangling ability of latent actions. To provide a more intuitive illustration, we randomly collect 1000 samples for each action from Habitat, Minecraft, and DMLab, and use UMAP (McInnes et al., 2018) for visualization. Figure 7 shows that the same actions, even from different environments, are clustered together, which validates the context-invariant property of our latent actions. Note that noise exists because the action inputs cannot be executed in certain states (e.g., cannot go ahead when an obstacle is in front). We also compare samples inferred by a model trained with a lower $\beta$. Although this results in more differentiable latent actions, it also reduces action overlap across environments thus sacrificing disentanglement ability. We therefore set $\beta$ as $2\times 10^{-4}$ by default.
Figure 7: UMAP of latent actions. Reducing the value of $\beta$ increases expressiveness but sacrifices disentanglement from context.
@@ 4 Conclusion
Acknowledgements
Shenyuan Gao and Jun Zhang were supported by the Hong Kong Research Grants Council under the NSFC/RGC Collaborative Research Scheme grant CRS_HKUST603/22.
@@ A.2 Data Mixture
Table 7: Data organization. Data sources, generation procedures, approximated frame counts, and mixture ratios for our training.
Category Data Source Automated # Frames Ratios
2D Video Game Gym Retro (Nichol et al., 2018) 1000M 49%
Procgen Benchmark (Cobbe et al., 2020) 144M 2%
@@ B.3 Sampling
B.4 Visual Planning on the Procgen Benchmark
@@
B.5 Visual Planning on the VP2 Benchmark
We follow the official protocol to evaluate all models on the VP2 benchmark. During adaptation, we use 5K given trajectories for Robosuite and 35K scripted trajectories with perturbations for RoboDesk for finetuning. Each world model is adapted for 1K steps with a batch size of 32 and a learning rate of $5\times 10^{-4}$. A cost below 0.05 is considered as success in Robosuite tabletop pushing tasks.
B.6 iVideoGPT Training Details
We resume from the official OpenX checkpoint of iVideoGPT and do not finetune its tokenizer. The model is designed to predict 15 future frames from an initial frame. After pretraining iVideoGPT and our action-aware variant on OpenX for 27K extra steps, we finetune each model with robot actions for 1K steps on BAIR robot pushing dataset. The resulting models are tested on 256 test videos.
@@ C.1 Action Transfer
Visualizations. We attach more action transfer results in a number of environments in Figures 12, 13, 14, 15 and 16. Each sample is generated by transferring a latent action sequence with a length of 20.
@@ Appendix E Selected Scenes for Visual Planning
To guarantee an effective evaluation, we conduct an exhaustive search using the ground truth environments to identify viable Procgen test scenes that can be completed in an acceptable number of steps by goal image matching. This results in a total of 120 scenes for our visual planning experiments in Sec. 3.2.2. In Figures 20, 17, 18 and 19, we illustrate the initial states of all selected scenes. The scenes from survival-oriented environments (e.g., BigFish and BossFight) are not considered in our evaluation.
Appendix F Related Work
F.1 World Models
Driven by advances in deep generative models (Peebles & Xie, 2023; Yang et al., 2025), world models have made encouraging strides recently. They aim to replicate the transitions of the world through generation (Sutton, 1991; Ha & Schmidhuber, 2018; Yang et al., 2024c; Lu et al., 2024, 2025; Agarwal et al., 2025), enabling agents to perform decision making entirely in imagination even when confronted with unseen situations (Kaiser et al., 2020; Micheli et al., 2023; Hafner et al., 2023; Du et al., 2023; Mazzaglia et al., 2024). This capability has proven beneficial for gaming agents (Kim et al., 2020; Alonso et al., 2024; Pearce et al., 2024; He et al., 2025), autonomous vehicles (Kim et al., 2021; Hu et al., 2023; Yang et al., 2024a; Gao et al., 2024; Wang et al., 2024; Bar et al., 2025; Hassan et al., 2025), and real robots (Seo et al., 2023; Wu et al., 2022; Ko et al., 2024; Du et al., 2024; Zhen et al., 2024; Zhou et al., 2024b; Wu et al., 2024; Zhu et al., 2024; Zhou et al., 2024a; Bu et al., 2024; Wang et al., 2025; Qi et al., 2025; Wu et al., 2025).
F.2 Latent Action from Videos
Humans can comprehend the notion of various actions by observing others to behave (Rizzolatti et al., 1996; Rybkin et al., 2019; Schmeckpeper et al., 2020; Yatim et al., 2024; Ling et al., 2025). To develop intelligent agents, it is intriguing to automatically extract similar action primitives from videos. Some approaches align human and robot manipulation videos in a shared space to derive meaningful action prototypes (Xu et al., 2023b). However, they require collecting semantically paired videos, which limits their scalability for most real-world tasks.
To avoid reliance on video pairs, some works extract implicit latent actions through fully unsupervised learning (Edwards et al., 2019; Menapace et al., 2021, 2022; Ye et al., 2023; Schmidt & Jiang, 2024; Zhang et al., 2024; Sun et al., 2024; Villar-Corrales & Behnke, 2025). However, they mainly focus on a single environment with limited complexity. While more recent works extend similar approaches to multiple datasets (Cui et al., 2024; Chen et al., 2024b, c; Ye et al., 2025; Bu et al., 2025; Ren et al., 2025), they mainly leverage latent actions as the objectives for behavior cloning. Consequently, they always adopt discrete actions to fit the policy networks (Kim et al., 2024), leading to significant ambiguity due to discretization (Nikulin et al., 2025). Thus, the potential of latent actions for adaptable world modeling remains underexplored.